### v1

@@ Abstract
In this paper, we experiment with novelty-based variants of OpenAI-ES, the NS-ES and NSR-ES algorithms, and evaluate their effectiveness in training complex, transformer-based architectures designed for the problem of reinforcement learning such as Decision Transformers. We also test if we can accelerate the novelty-based training of these larger models by seeding the training by a pretrained models. By this, we build on our previous work, where we tested the ability of evolution strategies – specifically the aforementioned OpenAI-ES – to train the Decision Transformer architecture. The results were mixed. NS-ES showed progress, but it would clearly need many more iterations for it to yield interesting results. NSR-ES, on the other hand, proved quite capable of being straightforwardly used on larger models, since its performance appears as similar between the feed-forward model and Decision Transformer, as it was for the OpenAI-ES in our previous work.
@@ I Introduction
Reinforcement learning is considered possibly the most difficult, yet also the most general, and in the future hopefully the most useful subfield of machine learning. Among many approaches to solving it [1], we can find those based on computing a gradient to optimize the objective, but also others that are derivative-free. Evolutionary algorithms [2] are one such class of general derivative-free optimization algorithms that can be used to solve this problem. Evolution strategies [3], which belong to this algorithmic family, have been proved to be a viable and competitive alternative to gradient approaches for the (deep) reinforcement learning [4]. Even though generally the gradient approaches have better sample utilization, the evolution strategies – just as many of their cousins from the family of evolutionary algorithms – are greatly parallelizable, which helps them overcome this limitation. Another benefit of using an evolutionary algorithm for reinforcement learning is that they have better exploration of possible solutions; therefore, agents trained using evolution strategies are usually more diverse than those trained by gradient-based algorithms.
@@
On a different note, the transformer architecture [10] has recently become the preferred solution in the field of neural networks and supervised learning for an ever-growing range of problems. In particular, there have been efforts to reinterpret reinforcement learning as a sequence modeling problem, utilizing the strengths of transformers to develop novel approaches for tackling such challenges. This has led to models like the Decision Transformer [11] and Trajectory Transformer [12]. Initially introduced as a model for offline reinforcement learning based on supervised sequence prediction, its authors also claim that the Decision Transformer performs effectively in traditional online reinforcement learning tasks as well.
In our previous work [13], we subjected the combination of the OpenAI-ES and the Decision Transformers to experiments testing the ability of derivative-free algorithms to train this more complicated and larger transformer architecture, compared to the simple feed-forward models that had been experimented with before. The evolution strategy proved to be a viable method to train the transformer in this setting. As a next step, we decided to test whether the novelty still provides enough training signals even for these larger models, and so we conducted experiments with NS-ES and NSR-ES testing their capability to train Decision Transformers.
@@ II-A Evolution Strategies
Evolution strategies are quite a successful family of black-box nature-inspired derivative-free optimization algorithms. They were introduced as a tool for solving high-dimensional continuous-valued problems [3]. The evolution strategies work with a population of real-valued vectors (called individuals). In each iteration (generation), they derive a new set of individuals by pertrubating (mutating) the original population; the new set is then evaluated with respect to a given objective function (fitness function), and a new generation is formed based on these new individuals taking into account their objective function value (fitness / fitness value).
To apply an evolution strategy as a reinforcement learning algorithm, a correspondence between the individuals in the algorithm and the reinforcement learning agents represented by a neural network is drawn using the network’s weights. The weights are vectorized and the resulting vectors of real numbers are then used as the individuals of the algorithm. The fitness of each individual is then defined as the average return of the corresponding agent over multiple episodes. Various evolution strategy algorithms have been proposed for this purpose [4].
@@
Nevertheless, evolution strategies also have many advantages. Numerous evolution strategy algorithms are highly parallelizable, and a significant portion of research in the field has been concentrated on this characteristic. As a result, we have algorithms that achieve a linear performance improvement as computing power increases [9]. Furthermore, since evolution strategies are derivative-free, they allow optimization not only of conventional smooth neural networks but also of models that include discrete subfunctions or other non-differentiable components. Another benefit is that compared to gradient-based algorithms, the evolution strategies have superior exploration.
This inherently good exploration can be furhter vastly increased by utilizing novelty [5, 6], which basically means searching for novel, previously unseen behavior, as compared to the classical fitness-based approach that seeks high-performing behaviors. The novelty can either completely replace the objective function, which yields us so-called novelty search algorithms [5, 6] – open-ended algorithms suitable when the rewards of an environment are not informative enough, when they are deceptive, or when they are hard to reasonably specify – or it can be used to complement the rewards, yielding us quality-diversity algorithms [7].
@@
In this paper, we will be working with NS-ES and NSR-ES [8], two variations of the OpenAI-ES [9] algorithm utilizing novelty. The first one is a pure novety search algorithm, the second one belongs to the quality-diversity algorithms. It will be beneficial for us to first understand the OpenAI-ES, and then extend this algorithm into the two tested in this paper.
OpenAI-ES is a representative of Natural evolution strategies [14]. It models the population as a probability distribution over the agent’s (neural network’s) parameters, specifically a Gaussian distribution. The mean of this distribution serves as the candidate solution to the given problem, while new offspring are generated by sampling from it each generation. These offspring are then evaluated, and their performance is used to update the distribution’s parameters (in our case, only the mean) to improve the expected fitness of future samples. This update follows an approximation of the natural gradient (whence the name of the algorithmic family). In our case, the natural gradient approximation is achieved by renormalizing (rescaling) the update based on uncertainty. In general, computing the natural gradient would involve inverting a so-called Fisher information matrix and applying it to the gradient estimate. However, as shown before in the literature [15], when deriving parameter updates from a Gaussian distribution with uniform variance across all parameters – just as we do – dividing by the variance (i.e., rescaling with respect to uncertainty) yields a similar effect and achieves a similar result. The algorithm is also designed for highly efficient parallelization, minimizing interprocess communication. For furhter details or a discussion of the design choices, we refer our readers to the original paper [9].
NS-ES and NSR-ES differ from OpenAI-ES just in a few things. First, they keep a metapopulation of several distributions (represented by their means) serving as distinct populations. Only the behaviors of distribution means are added to the archive. A member of the metapopulation that is to be improved in a given iteration is chosen proportionally to its current novelty. In NS-ES, wherever the fitness would be used, the novelty (computed with respect to the current state of the behavior archive) is used. As for the NSR-ES, the fitness is combined with the novelty by averaging the two.
@@ II-B Transformers
Transformers are currently the state-of-the-art sequence-to-sequence neural architecture utilized for numerous tasks of supervised learning [10]. As a rule of thumb, they appear to possess strong generalization capabilities; the greater the larger the model employed. However, achieving these impressive results also demands a substantial amount of training data.
The most important component, to which the transformers owe their success, is a self-attention layer, used repeatedly throughout the network. For each input sequence element, the self-attention constructs a "key", a "query", and a "value". Next, an $i$-th output element is obtained as a linear combination of all values, with each value weighted in proportion to the product of the query at the given ($i$-th) position and the key associated with the value. Thus, this integrates information from the entire input sequence to generate each individual element of the output, as again expressed in the following equation.
@@
We can apply a mask and for each element concealing the portion of the input sequence which follows it. This ensures that only preceding information is used to derive the output element. This technique is known as causal masking, and a transformer which implements it is referred to as a causal transformer.
@@
Now, let us examine the proposed architecture itself. It consists of a causal transformer; embeddings for returns-to-go, observations (states of the environment), and actions; position encoder; and a linear decoder to transform the output of the transformer into actions, as shown in Figure 1.
@@ III Experiments
Figure 2: OpenAI-ES, NS-ES and NSR-ES used on a simple feed-forward model (FF), a Decision Transformer (DT), and a pretrained Decision Transformer (Pre) for a MuJoCo Humanoid simulation in ten runs of each of the experiments. The data from the ten runs are aggregated together and the mean values and percentile intervals are shown. Figures 2(a), 2(c) and 2(e) show solution evaluation fitness, with the width of the percentile interval being 100 %, and Figures 2(b), 2(d) and 2(f) show runtimes of population evaluation episodes with standard 97,5 % interval.
In our previous work [13], we tested the ability of OpenAI-ES to train agents with their policy constituted by a Decision Transformer. We then also proposed a method of aiding this evolution strategy in training large models using first a pretraining of the large model in a form of a behavior cloning towards some smaller, easily trained, yet possibly weaker model. The results were promising. The algorithm was mostly capable of training Decision Transformers, even without the pretraining.
A logical next step is to test whether the novelty described in the previous section provides us enough information to train these bigger models. Therefore, we extended our previous implementation of OpenAI-ES into implementations of NS-ES and NSR-ES and proceeded to test these novelty search and quality-diversity algorithms in the MuJoCo [16] Humanoid environment using OpenAI Gym [17]. For details of the implementation, we refer our reader to the original papers for OpenAI-ES [9] and for NS-ES / NSR-ES [8] regarding overall details, and to our previous paper [13] regarding our tweaks of the implementation and their justification, as our current code simply extends the previous implementation into a novelty-utilizing form.
For both inspected algorithms, we first conducted a replication experiment of the original paper and a correctness check of our implementation using a classical feed-forward network, which then also served as a baseline for follow-up experiments. We then proceeded with experiments on Decision Transformers and concluded by testing whether the pretraining helps to accelerate the novelty-based training of these larger models.
For all the experiments, we conducted ten runs of the training. For each run of each of the experiments, 300 workers were used utilizing 301 CPU cores (with one being a master handling synchronization, evaluation, and saving the agent). The desired returns passed to all the Decision Transformer models at the beginning of each episode were 7000.
Unless stated otherwise in the individual experiment descriptions, all the remaining hyperparameter values can be found as default values in our codebase. [^1] Our code and the data collected during the conducted experiments can be found on GitHub repository on the following link: https://github.com/Mafi412/Novelty-based-Evolution-Strategies-and-Decision-Transformers
As a baseline for the novelty-based algorithms tested in this work, we present the results that the purely objective-based algorithm, OpenAI-ES, was able to achieve during the experiments from the previous paper [13]. These can be seen in Figures 2(a) and 2(b).
@@ III-A NS-ES
Figure 3: Comparison of the average distances that the resulting agents trained by the examined algorithms were able to travel from their starting positions in the Humanoid environment. Algorithms in question are objective-based OpenAI-ES; novelty search NS-ES; quality-diversity NSR-ES; and Random standing for randomly initialized agents without any further training. When an algorithm works with a metapopulation, the agent with the best average distance in the final metapopulation was chosen as a solution for this plot. Finally, for each algorithm in question (except for Random), results of all three types of experiments are plotted – training of a feed-forward model (FF) and training of a Decision Transformer without and with the pretraining (DT and DT - Pretrained). For Random, only a feed-forward model was used.
@@
We start with the novelty search algorithm, NS-ES. We started with a replication experiment, where we used our implementation of this algorithm to train the feed-forward model used in the original paper [8]. This also gave us a good baseline for our experiments with Decision Transformers. We then used the same algorithm to train the Decision Transformer from scratch; the only difference was that because the transformer is almost five times larger, we quadrupled the size of the population the algorithm works with. The last conducted experiment for this section was testing if seeding the training with a pretrained Decision Transformer accelerates the process, as the training using just a novelty signal is usually less efficient than following an objective; hence it might prove beneficial to be able to speed it up, more so when training larger models. In accordance with our previous work [13], no virtual batch normalization was used and values of learning rate and noise deviation hyperparameters were both reduced to $0.01$ when utilizing the pretraining. The results of these experiments can be seen in Figures 2(c) and 2(d).
@@
Since our objective shifts when using novelty from "teaching the agent to walk straight forward as efficiently as possible" towards simply "teaching the agent to walk", the fitness is no longer the best representation of agent’s progress. Much more informative are the runtimes in this case, which tell us how long was the agent able to stay on its feet, and hence Figure 2(d) is more important to us now. There, we can observe that the progression in training the Decision Transformer when not using the pretraining does indeed occur, but much later than when training the feed-forward model. That is in stark contrast to the purely objective-based case shown in Figures 2(a) and 2(b), where the difference between the feed-forward and Decision Transformer cases is not so significant. As for the pretraining, it only seems to hurt the training in this case, as no progress can be seen in Figure 2(d). Still, in Figure 2(a), we can see that some progress is being made even with respect to the fitness for the Decision Transformers, at least without the pretraining.
@@
In Figure 2(d), we can see a dent for the feed-forward data when, after reaching a point where a part of the population generated in each generation is able to stay on their feet, this progress is reverted for a few iterations before the generated population regains this ability. This is caused solely by switching between various members of the metapopulation based on their current novelty throughout the training, which is a feature of the algorithm.
@@
Of course, one might ask about the final performance of our trained agents with respect to our above-declared objective: How far are the agents able to walk after the training? This can be seen in Figure 3. We can clearly see that even though the novelty signal is capable of somewhat training the larger models, it would need much more time to achieve similar results as with the smaller feed-forward models. And we can see that the pretrained Decision Transformers at least fared better after the novelty search training than the random agents, but this has to be viewed in context of how long the training was and that the pretrained agents were better before the training.

### v2

@@ Abstract
In this paper, we experiment with novelty-based variants of OpenAI-ES, the NS-ES and NSR-ES algorithms, and evaluate their effectiveness in training complex transformer-based architectures designed for the problem of reinforcement learning, such as Decision Transformers. We also test whether we can accelerate the novelty-based training of these larger models by seeding the training with a pretrained model. The experimental results were mixed. NS-ES showed progress, but it would clearly need many more iterations for it to yield interesting agents. NSR-ES, on the other hand, proved quite capable of being straightforwardly used on larger models, since its performance appears as similar between the feed-forward model and Decision Transformer, as it is for the baseline objective-based OpenAI-ES.
@@ I Introduction
Reinforcement learning is considered possibly the most difficult and in the future hopefully the most useful subfield of machine learning, since it mimics how humans and animals learn [1]. Among many approaches to solving it [1], we can find those based on computing a gradient to optimize the objective, but also others that are derivative-free. Evolutionary algorithms [2] are one such class of general derivative-free optimization algorithms that can be used to solve this problem. Evolution strategies [3], which belong to this algorithmic family, have been proved to be a viable and competitive alternative to gradient approaches for the (deep) reinforcement learning [4]. Even though generally, the gradient approaches have better sample utilization, the evolution strategies – just as many of their cousins from the family of evolutionary algorithms – are greatly parallelizable, which helps them overcome this limitation. Another benefit of using an evolutionary algorithm for reinforcement learning is that they have better exploration of possible solutions; therefore, agents trained using evolution strategies are usually more diverse than those trained by gradient-based algorithms [5, 6].
@@
On a different note, a transformer architecture [12] has recently become the preferred solution in the field of neural networks and supervised learning for an ever-growing range of problems, be it its original task of language processing [12] or an image processing task [13]. In particular, there have been efforts to reinterpret a reinforcement learning as a sequence modeling problem, utilizing the strengths of transformers to develop novel approaches for tackling such challenges. This has led to models like the Decision Transformer [14] and Trajectory Transformer [15]. Initially introduced as a model for offline reinforcement learning based on supervised sequence prediction, its authors also claim that the Decision Transformer performs effectively in traditional online reinforcement learning tasks as well.
In our previous work [16], we subjected the combination of the OpenAI-ES and the Decision Transformers to experiments testing the ability of derivative-free algorithms to train this more complicated and larger transformer architecture, compared to the simple feed-forward models that had been experimented with before. The evolution strategy proved to be a viable method to train the transformer in this setting. Still, the OpenAI-ES employs a clear fitness guidance for training, so as a next step, we decided to test whether the novelty still provides enough training signals even for these larger models, and hence we conducted experiments with NS-ES and NSR-ES testing their capability to train Decision Transformers.
The contribution of this paper thus consists in demonstrating the ability of the aforementioned novelty-based approaches to train agents whose policies are based on a transformer architecture in a reinforcement learning setting. [^1] Our code, together with the data from all the conducted experiments can be found on GitHub repository on the following link: https://github.com/Mafi412/Novelty-based-Evolution-Strategies-and-Decision-Transformers
@@ II-A Evolution Strategies
Evolution strategies are quite a successful family of black-box nature-inspired derivative-free optimization algorithms. They were introduced as a tool for solving high-dimensional continuous-valued problems [3]. The evolution strategies work with a population of real-valued vectors (called individuals). In each iteration (generation), they derive a new set of individuals by perturbating (mutating) the original population; the new set is then evaluated with respect to a given objective function (fitness function), and a new generation is formed based on these new individuals taking into account their objective function value (fitness or fitness value).
To apply an evolution strategy as a reinforcement learning algorithm, a correspondence between the individuals in the algorithm and the reinforcement learning agents represented by a neural network is drawn using the network’s weights. The weights are vectorized and the resulting vectors of real numbers are then used as the individuals for the algorithm. The fitness of each individual is then defined as the average cumulative reward – called return – of the corresponding agent over multiple episodes. Various evolution strategy algorithms have been proposed for this purpose [4].
@@
Nevertheless, evolution strategies also have many advantages. Numerous evolution strategy algorithms are highly parallelizable, and a significant portion of the research in the field has been concentrated on this characteristic. As a result, we have algorithms that achieve a linear performance improvement as the computing power increases [11]. Furthermore, since evolution strategies are derivative-free, they allow optimization not only of conventional smooth neural networks but also of models that include discrete subfunctions or other non-differentiable components. Another benefit is that compared to gradient-based algorithms, the evolution strategies have superior exploration [5, 6].
This inherently good exploration can be further vastly increased by utilizing novelty [7, 8], which basically means searching for novel, previously unseen behavior, as compared to the classical fitness-based approach that seeks high-performing behaviors. The novelty can either completely replace the objective function, which yields us so-called novelty search algorithms [7, 8] – open-ended algorithms suitable when the rewards of an environment are not informative enough, when they are deceptive, or when they are hard to reasonably specify – or it can be used to complement the rewards, yielding us quality-diversity algorithms [9].
@@
In this paper, we will be working with NS-ES and NSR-ES [10], two variations of the OpenAI-ES [11] algorithm utilizing novelty. The first one is a pure novelty search algorithm; the second one belongs to the quality-diversity algorithms. It will be beneficial for us to first understand the OpenAI-ES and then extend this algorithm into the two tested in this paper.
OpenAI-ES is a representative of Natural evolution strategies [17]. It models the population as a probability distribution over the agent’s (neural network’s) parameters, specifically a Gaussian distribution. The mean of this distribution serves as the candidate solution to the given problem, while new offspring are generated by sampling from it each generation. These offspring are then evaluated, and their performance is used to update the distribution’s parameters (in our case, only the mean) to improve the expected fitness of future samples. This update follows an approximation of the natural gradient (whence the name of the algorithmic family). In our case, the natural gradient approximation is achieved by renormalizing (rescaling) the update based on uncertainty. In general, computing the natural gradient would involve inverting a so-called Fisher information matrix and applying it to the gradient estimate. However, as shown before in the literature [18], when deriving parameter updates from a Gaussian distribution with uniform variance across all parameters – just as we do – dividing by the variance (i.e., rescaling with respect to uncertainty) yields a similar effect and achieves a similar result. The algorithm is also designed for highly efficient parallelization, minimizing interprocess communication. For further details or a discussion of the design choices, we refer our readers to the original paper [11].
NS-ES and NSR-ES differ from OpenAI-ES just in a few things. First, they maintain a metapopulation of several distributions (represented by their means) serving as distinct populations. Only the behaviors of the distribution means are added to the archive. A member of the metapopulation to be improved in a given iteration is chosen proportionally to its current novelty. In NS-ES, wherever the fitness would be used, the novelty (computed with respect to the current state of the behavior archive) is used. As for the NSR-ES, the fitness is combined with the novelty by averaging the two.
@@ II-B Transformers
Transformers are currently the state-of-the-art sequence-to-sequence neural architecture utilized for numerous tasks of supervised learning [12, 13]. As a rule of thumb, they appear to possess strong generalization capabilities; the greater the larger the model employed. However, achieving these impressive results also demands a substantial amount of training data.
The most important component, to which the transformers owe their success, is a self-attention layer, which is used repeatedly throughout the network. For each input sequence element, the self-attention constructs a "key", a "query", and a "value". Next, an $i$-th output element is obtained as a linear combination of all values, with each value weighted in proportion to the product of the query at the given ($i$-th) position and the key associated with the value. Thus, this integrates information from the entire input sequence to generate each individual element of the output, as again expressed in the following equation.
@@
We can apply a mask and for each element conceal the portion of the input sequence which follows it. This ensures that only preceding information is used to derive the output element. This technique is known as a causal masking, and a transformer which implements it is referred to as a causal transformer.
@@
Now, let us examine the proposed architecture itself. It consists of a causal transformer; embeddings for returns-to-go, observations (states of the environment), and actions; position encoder; and a linear decoder to transform the output of the transformer into actions, as shown in Fig. 1.
@@ III Experiments
In our previous work [16], we tested the ability of OpenAI-ES to train agents with their policy constituted by a Decision Transformer. We also proposed a method for aiding this evolution strategy in training large models using first a pretraining of the large model in a form of a behavior cloning towards some smaller, easily trained, yet possibly weaker model. The results were promising. The algorithm was mostly capable of training Decision Transformers, even without the pretraining.
A logical next step is to test whether the novelty described in the previous section provides us with enough information to train these larger models. Therefore, we extended our implementation of OpenAI-ES into implementations of NS-ES and NSR-ES and proceeded to test these novelty search and quality-diversity algorithms in the MuJoCo [19] Humanoid environment using OpenAI Gym [20]. For details of the implementation, we refer our reader to the original papers for OpenAI-ES [11] and for NS-ES / NSR-ES [10] regarding overall details, and to our previous paper [16] regarding a few tweaks in the implementation and their justifications, as our current code simply extends the previous implementation into a novelty-utilizing form.
Because training a larger model – which the Decision Transformer is – is time-intensive, we chose to test the performance of the algorithms only in the Humanoid environment. It should be a suitable representative of MuJoCo environments, as it is the most complex and challenging among the standard ones.
As mentioned in Section II-A, utilizing novelty requires a behavior characteristic. We use the standard one for MuJoCo environments: the behavior characteristic of an agent is defined as its final position in the environment – that is, the $x$ and $y$ coordinates at which the agent concludes the episode. The distance between two such characteristics is defined as a classical Euclidean distance between the two points.
For both the inspected algorithms, we first conducted a replication experiment of the original paper [10] and a correctness check of our implementation using a classical feed-forward network, which then also served as a baseline for follow-up experiments. We then proceeded with experiments on Decision Transformers and concluded by testing whether the pretraining helps to accelerate the novelty-based training of these larger models.
For all the experiments, we conducted ten runs of the training. For each run of each of the experiments, 300 workers were used utilizing 301 CPU cores (with one being a master handling synchronization, evaluation, and saving the agent). The desired returns passed to all the Decision Transformer models at the beginning of each episode were 7000. However, it is important to note that in the original Decision Transformer paper [14], rewards and returns-to-go are scaled down by a factor of 1000 in MuJoCo environments. To stay consistent with this approach, we have adopted the same rescaling. All figures showing fitness progression during the evolution training in the Humanoid environment are based on these scaled returns, as this is the format in which the values are logged during training. To retrieve the actual fitness values as produced by the original Humanoid environment, one simply needs to multiply the reported numbers by 1000.
Unless stated otherwise in the individual experiment descriptions, all the remaining hyperparameter values can be found as default values of the arguments in our codebase.
As a baseline for the novelty-based algorithms tested in this work, we present the results that the purely objective-based algorithm, OpenAI-ES, was able to achieve. These can be seen in Fig. 2(a) and 2(b). Even though there are no purely gradient-based reinforcement learning algorithms utilizing novelty, it might still be interesting to compare the evolution-based approaches shown here with some gradient-based algorithm. Therefore, we show the best final average performance achieved by training a Decision Transformer using TD3 algorithm during roughly the same wall-clock time as the OpenAI-ES. All these objective-based results were adopted from an earlier paper [16].
@@
Figure 2: OpenAI-ES, NS-ES and NSR-ES used on a simple feed-forward model ($F$), a Decision Transformer (DT), and a pretrained Decision Transformer (Pre) for a MuJoCo Humanoid simulation in ten runs of each of the experiments. Fig. 2(a), 2(c), and 2(e) show solution evaluation fitness and Fig. 2(b), 2(d), and 2(f) show runtimes of population evaluation episodes. The data from the ten runs are aggregated together: For the evaluation results, the median and quartiles are shown, whereas for the runtimes (where we have a lot more values per iteration), the mean values and percentile intervals (97,5 %) are displayed. The TD3 horizontal line in Fig. 2(a) shows the best final average performance of the TD3 algorithm used on a Decision Transformer. (In Fig. 2(a), the ‘Pre’ line ends early because, for OpenAI-ES, the experiments with pretrained agents were shorter than those without pretraining and used only 100 iterations, compared to 200 for the others.)
@@ III-A NS-ES
We start with the novelty search algorithm, NS-ES. We started with a replication experiment, where we used our implementation of this algorithm to train the feed-forward model used in the original paper [10]. This also gave us a good baseline for our experiments with Decision Transformers. We then used the same algorithm to train the Decision Transformer from scratch; the only difference was that because the transformer is almost five times larger, we quadrupled the size of the population the algorithm works with. (This, of course, increases the number of function evaluations – episodes of interactions with the environment – by a factor of four as well.) The last conducted experiment for this section was testing if seeding the training with a pretrained Decision Transformer accelerates the process, as the training using just a novelty signal is usually less efficient than following an objective; hence it might prove beneficial to be able to speed it up, more so when training larger models. In accordance with previous research on utilizing OpenAI-ES with transformers, no virtual batch normalization was used and values of learning rate and noise deviation hyperparameters were both reduced to $0.01$ when utilizing the pretraining. The results of these experiments can be seen in Fig. 2(c) and 2(d).
@@
Since our objective shifts when using novelty from “teaching the agent to walk straight forward as efficiently as possible” toward simply “teaching the agent to walk”, the fitness is no longer the best representation of the agent’s progress. (Still, it tells us how are we faring with respect to the original objective.) Much more informative are the runtimes in this case, which tell us how long was the agent able to stay on its feet, and hence Fig. 2(d) is more important to us now. There, we can observe that the progression in training the Decision Transformer when not using the pretraining does indeed occur, but much later than when training the feed-forward model. That is in stark contrast to the purely objective-based case shown in Fig. 2(a) and 2(b), where the difference between the feed-forward and Decision Transformer cases is not so significant. As for the pretraining, it only seems to hurt the training in this case, as no progress can be seen in Fig. 2(d). Still, in Fig. 2(c), we can see that some progress is being made even with respect to the fitness for the Decision Transformers, at least without the pretraining.
@@
In Fig. 2(d), we can see a dent for the feed-forward data when, after reaching a point where a part of the population generated in each generation is able to stay on their feet, this progress is reverted for a few iterations before the generated population regains this ability. This is caused solely by switching between various members of the metapopulation based on their current novelty throughout the training, which is a feature of the algorithm.
@@
Of course, one might ask about the final performance of our trained agents with respect to our above-declared objective: How far are the agents able to walk after the training? This can be seen in Fig. 3. We can clearly see that even though the novelty signal is capable of somewhat training the larger models, it would need much more time to achieve similar results as with the smaller feed-forward models. And we can see that the pretrained Decision Transformers at least fared better after the novelty search training than the random agents, but this has to be viewed in context of how long the training was and that the pretrained agents were better before the training.
Figure 3: Comparison of the average distances that the resulting agents trained by the examined algorithms were able to travel from their starting positions in the Humanoid environment. Algorithms in question are objective-based OpenAI-ES; novelty search NS-ES; quality-diversity NSR-ES; and Random standing for randomly initialized agents without any further training. When an algorithm works with a metapopulation, the agent with the best average distance in the final metapopulation was chosen as a solution for this plot. Finally, for each algorithm in question (except for Random), results of all three types of experiments are plotted – training of a feed-forward model ($F$) and training of a Decision Transformer without and with the pretraining (DT and DT - Pretrained, respectively). For Random, only the feed-forward model was used.